# A dimensional mass transference principle from balls to open sets and applications to dynamical Diophantine approximation

### v1

@@ 2. Preliminary
Let $A$ be a Borel subset of $\mathbb{R}^{d}$. If $A$ supports a Borel probability measure $\lambda$ that satisfies
@@
for some constant $0<c<\infty$, and for every ball $x\in\mathbb{R}^{d}$ and $r>0$, then
$\mathcal{H}_{\infty}^{s}(A)\geq 1/c.$
Lemma 2.2 (Frostman’s lemma [33, Theorem 8.8]).
Let $A$ be a Borel subset of $\mathbb{R}^{d}$. If $\mathcal{H}_{\infty}^{s}(A)>0$, then there exists a probability measure $\lambda$ supported on $A$ such that for any $x\in\mathbb{R}^{d}$ and $r>0$,
$\lambda(B(x,r))\ll\frac{r^{s}}{\mathcal{H}_{\infty}^{s}(A)},$
@@
Theorem 2.3 offers a relatively simple criterion for verifying that a $\limsup$ set has the large intersection property.
Theorem 2.3 ([20, Corollary 2.6]).
@@
$\limsup_{n\to\infty}\mathcal{H}_{\infty}^{t}(E_{n}\cap B)>c\mu(B)$
@@
Theorem 2.4 (Besicovitch covering Theorem [33, Theorem 2.7]).
There is a positive integer $Q_{d}$ depending only on the dimension $d$ with the following property. Let $A\subset\mathbb{R}^{d}$ be a bounded set, and let $\mathcal{B}$ be a family of balls such that each point of $A$ is the centre of some ball of $\mathcal{B}$. There are families $\mathcal{B}_{1},\dots,\mathcal{B}_{Q_{d}}\subset\mathcal{B}$ covering $A$ such that each $\mathcal{B}_{k}$ is disjoint, that is,
$A\subset\bigcup_{1\leq k\leq Q_{d}}\bigcup_{B\in\mathcal{B}_{k}}B$
@@
Lemma 2.5 ([7, Lemma 5]).
Let $\{B(x_{n},r_{n})\}$ be a sequence of balls in $X\subset\mathbb{R}^{d}$ with $r_{n}\to 0$ as $n\to\infty$. Then, for any ball $B$ in $X$, there exists a finite collection
@@
    there exists a constant $c>0$ independent of $B$ such that
    $\mu\bigg{(}\bigcup_{B(x_{n},r_{n})\in K_{B}}B(x_{n},r_{n})\bigg{)}\geq c\mu(B).$
Here, we highlight the difference between Besicovitch covering theorem and Lemma 2.5. Besicovitch covering theorem applies to arbitrary measures but is restricted to Euclidean spaces, whereas Lemma 2.5 can be extended to general metric spaces but requires the measure to be doubling. In the sequel, when it is necessary to extract a disjoint subcollection from a sequence of shrinking balls, we will use Lemma 2.5 for the ambient measure $\mu$, and the Besicovitch covering theorem for the reference measure $\nu$.
Note that as the collection $\{F_{n}\}$ of closed sets in Theorem 1.4 are not necessarily balls, the following variant of Lemma 2.5 is needed.
Corollary 2.6.
Let $\mathcal{F}=\{F_{n}\}$ be as given in Theorem 1.4. That is, each $F_{n}\in\mathcal{F}$ satisfies Definition 1.2 (1) and
$\mu\Big{(}\limsup_{n\to\infty}F_{n}\Big{)}=1.$
@@
    If $F_{n}$ and $F_{m}$ are distinct then $\mathrm{dist}(F_{n},F_{m})\geq\max\{|F_{n}|,|F_{n^{\prime}}|\}$, where the distance between sets is defined by
@@
    there exists a constant $c^{\prime}>0$ independent of $B$ such that
    $\mu\bigg{(}\bigcup_{F_{n}\in\mathcal{F}_{B}}F_{n}\bigg{)}\geq c^{\prime}\mu(B).$
@@
Recall from Remark 1 (b) that there exists an absolute constant $c\geq 1$ such that for any $F_{n}$,
@@
where $x_{n}\in F_{n}$. Clearly,
$\mu\Big{(}\limsup_{n\to\infty}B(x_{n},c|F_{n}|)\Big{)}=1$
Applying Lemma 2.5 to the sequence $\{B(x_{n},c|F_{n}|)\}$, we obtain a finite collection $K_{B}$ satisfying items (1)–(3) in Lemma 2.5. Let
@@
By (2.1), items (1) and (3) in corollary follows immediately from those in Lemma 2.5.
For item (2), suppose that $F_{n},F_{m}\in\mathcal{F}_{B}$ are distinct. Then, by definition, the same is true for $B(x_{n},c|F_{n}|),B(x_{m},c|F_{m}|)\in K_{B}$. It follows that
@@
$\mathrm{dist}(F_{n},F_{m})\geq\mathrm{dist}\big{(}B(x_{n},c|F_{n}|),B(x_{m},c|F_{m}|)\big{)}\geq\max\{|F_{n}|,|F_{m}|\}.\qed$
@@
Let $\{E_{n}\}$ be as given in Theorem 1.4. In this section, our goal is to establish that the corresponding $\limsup$ set has the large intersection property. Specifically, we will show that for any $0<t<s$,
$\limsup_{\ell\to\infty}\mathcal{H}_{\infty}^{t}\bigg{(}\bigcup_{k=\ell}^{\infty}E_{k}\cap B\bigg{)}\gg\mu(B)\quad\text{holds for all ball $B\subset X$},$
@@ 3.1. Construction of a subset of $\bigcup_{k=\ell}^{\infty}E_{k}\cap B$
For the given ball $B$, let $\mathcal{F}_{B}\subset\mathcal{F}$ be the finite subcollection of balls obtained from Corollary 2.6. Recall that for any $F_{i}\subset\mathcal{F}$,
@@
where $f_{i}:F_{i}\to X$ is a bijection. Let $F_{i}\in\mathcal{F}_{B}$. For any $x\in F_{i}$ and $0<\rho<|F_{i}|$, it follows from Definition 1.2 that $f_{i}$ is injective on $B(x,\rho)\cap F_{i}$. Moreover, we estimate the measure of a small ball intersected with $F_{i}$ as follows:
@@
(3.4) $\leq C\nu(F_{i})\nu\big{(}B(f_{i}(x),C\rho/|F_{i}|)\big{)},$
where the last inequality follows from (3.3) and item (1) of Remark 1.
@@
Then, for any $0<\rho<|F_{i}|/(CN)$ (or equivalently $0<C\rho/|F_{i}|<1/N$), applying (3.4) and the above inequality yields
$\begin{split}\nu(B(x,\rho)\cap F_{i})&\leq C\nu(F_{i})\nu\bigg{(}B\bigg{(}f_{i}(x),\frac{C\rho}{|F_{i}|}\bigg{)}\bigg{)}\leq C\nu(F_{i})\cdot\bigg{(}\frac{C\rho}{|F_{i}|}\bigg{)}^{\operatorname{\overline{\dim}_{H}}\nu-\varepsilon}\\ &\leq C^{d+1}\nu(F_{i})\cdot\bigg{(}\frac{\rho}{|F_{i}|}\bigg{)}^{\operatorname{\overline{\dim}_{H}}\nu-\varepsilon}.\end{split}$
@@
Since $\nu(\limsup B(x_{n},r_{n}))=1$, we have $f_{i}^{-1}(E_{\nu}^{N,\varepsilon})\subset E_{\nu,F_{i}}^{N,\varepsilon}$ except for a set of zero $\nu$-measure. Therefore,
$\nu(E_{\nu,F_{i}}^{N,\varepsilon})\geq\nu\big{(}f_{i}^{-1}(E_{\nu}^{N,\varepsilon})\big{)}=\nu(F_{i})\nu^{(i)}(E_{\nu}^{N,\varepsilon})\geq\frac{\nu(F_{i})\nu(E_{\nu}^{N,\varepsilon})}{C}\geq\frac{\gamma_{\varepsilon}\nu(F_{i})}{2C},$
where the last inequality follows from (3.2).
@@
The above inclusion $B(x_{n_{z}},r_{n_{z}})\subset B$ is possible since $E_{\nu,F_{i}}^{N,\varepsilon}\subset F_{i}\subset B$ and $B$ is open. Set $L_{n_{z}}:=B(z,5r_{n_{z}})$. Then, we have
@@
Since $\nu(E_{\nu,F_{i}}^{N,\varepsilon})\geq\gamma_{\varepsilon}\nu(F_{i})/(2C)$, there exists some $1\leq k_{i}\leq Q_{d}$ such that the corresponding collection $\mathcal{B}_{k_{i}}(F_{i})$ satisfies
$\nu\bigg{(}\bigcup_{L_{n_{z}}\in\mathcal{B}_{k_{i}}(F_{i})}L_{n_{z}}\bigg{)}\geq\frac{\nu(E_{\nu,F_{i}}^{N,\varepsilon})}{Q_{d}}\geq\frac{\gamma_{\varepsilon}\nu(F_{i})}{2CQ_{d}}.$
By the disjointness of the balls in $\mathcal{B}_{k_{i}}(F_{i})$, we may further extract a finite subcollection $\mathcal{B}(F_{i})\subset\mathcal{B}_{k_{i}}(F_{i})$ such that
@@
(1) It follows from immediately from Corollary 2.6.

### v2

@@ 2. Preliminary
Let $E$ be a Borel subset of $\mathbb{R}^{d}$. Suppose that $E$ supports a Borel probability measure $\lambda$ that satisfies
@@
for all $x\in\mathbb{R}^{d}$ and $r>0$, where $0<c<\infty$ is a constant. Then
$\mathcal{H}_{\infty}^{s}(E)\geq 1/c.$
Lemma 2.2 (Frostman’s lemma [36, Theorem 8.8]).
Let $E$ be a Borel subset of $\mathbb{R}^{d}$. If $\mathcal{H}_{\infty}^{s}(E)>0$, then there exists a probability measure $\lambda$ supported on $E$ such that for any $x\in\mathbb{R}^{d}$ and $r>0$,
$\lambda(B(x,r))\ll\frac{r^{s}}{\mathcal{H}_{\infty}^{s}(E)},$
@@
Theorem 2.3 below offers a relatively simple criterion for verifying that a $\limsup$ set has the large intersection property.
Theorem 2.3 ([22, Corollary 2.6]).
@@
$\limsup_{n\to\infty}\mathcal{H}_{\infty}^{t}(E_{n}\cap B)>c\mu(B)\asymp|B|^{\delta}$
@@
Remark 6.
Although the Falconer’s original definition of large intersection property (see Definition 1.3) requires that
$\mathcal{H}_{\infty}^{t}(E\cap B)\gg\mathcal{H}_{\infty}^{t}(B)\asymp|B|^{t}\quad\text{holds for any ball $B\subset X$},$
it was observed by the author in [22] that this condition can be significantly weakened, as stated in the above theorem. It is also worth noting that several alternative definitions of large intersection properties appear in that paper, some of which may be used to derive Hausdorff measures. However, we will not pursue these here, since our focus is on Hausdorff dimension.
@@
Theorem 2.4 (Besicovitch covering Theorem [36, Theorem 2.7]).
There is a positive integer $Q_{d}$ depending only on the dimension $d$ with the following property. Let $E\subset\mathbb{R}^{d}$ be a bounded set, and let $\mathcal{B}$ be a family of balls such that each point of $E$ is the centre of some ball of $\mathcal{B}$. There are families $\mathcal{B}_{1},\dots,\mathcal{B}_{Q_{d}}\subset\mathcal{B}$ covering $E$ such that each $\mathcal{B}_{k}$ is disjoint, that is,
$E\subset\bigcup_{1\leq k\leq Q_{d}}\bigcup_{B\in\mathcal{B}_{k}}B$
@@
Lemma 2.5 ([1, Lemma 5]).
Let $\{B(x_{n},r_{n})\}$ be a sequence of balls in $X\subset\mathbb{R}^{d}$ such that
$\mu\bigg(\limsup_{n\to\infty}B(x_{n},r_{n})\bigg)=1\mathrm{\quad and\quad}\text{$r_{n}\to 0$ as $n\to\infty$.}$
Then, for any ball $B$ in $X$, there exists a finite collection
@@
    $\mu\big(\bigcup_{B(x_{n},r_{n})\in K_{B}}B(x_{n},r_{n})\big)\gg\mu(B),$ where the implied constant does not depend on $B$.
Here, we highlight the difference between Besicovitch covering theorem and Lemma 2.5. Besicovitch covering theorem applies to arbitrary Borel measures but is restricted to Euclidean spaces, whereas Lemma 2.5 can be extended to general metric spaces but requires the measure to be doubling. In the sequel, when it is necessary to extract a disjoint subcollection from a sequence of shrinking balls, we will use Lemma 2.5 for the ambient measure $\mu$, and the Besicovitch covering theorem for the reference measure $\nu$.
Note that as the collection $\{F_{n}\}$ of closed sets in Theorem 1.4 are not necessarily balls, the following variant of Lemma 2.5 is needed. Before moving on, we give some simple properties of the map $f_{n}$ and the set $F_{n}$ defined in Definition 1.2 that will be used soon.
Proposition 2.6.
Let $F$ be a closed subset of $X$ and $f:F\to X$ be a bijection. Suppose that there exists a constant $C\geq 1$ such that
(2.1) $C^{-1}\dfrac{|x-y|}{|F|}\leq|f(x)-f(y)|\leq C\dfrac{|x-y|}{|F|}\quad\text{for all $x,y\in F$},$
then the following properties hold:
  1. (1)
    For any $x\in F$ and $0<r<|F|$,
    $B(f(x),C^{-1}r/|F|)\subset f(B(x,r)\cap F)\subset B(f(x),Cr/|F|).$
  2. (2)
    there exists a constant $c>0$ depends on $X$, $C$ and $\mu$ only such that
    $F\subset B(x,c|F|)\mathrm{\quad and\quad}c^{-1}|F|^{\delta}\leq\mu(F)\leq c|F|^{\delta},$
    where $x\in F$.
Proof.
(1) Let $x\in F$ and $0<r<|F|$. For any $y\in B(x,r)\cap F$, we have
[…]
and therefore $f(B(x,r)\cap F)\subset B(f(x),Cr/|F|)$.
For the left inclusion in item (1), take an arbitrary point $z$ in $B(f(x),C^{-1}r/|F|)$. Since $f$ is a bijection, there exists $y\in F$ such that $f(y)=z$. It follows that
[…]
Therefore, $y\in B(x,r)\cap F$, and consequently, $z=f(y)\in f(B(x,r)\cap F)$.
(2) Suppose, for the sake of contradiction, that $F\nsubseteq B(x,2C\,|X|\cdot|F|).$ Then, for any $y\in F\setminus B(x,2C|X|\cdot|F|)$, we would have
[…]
which is a contradiction. Hence,
$F\subset B(x,2C\,|X|\cdot|F|).$
It then follows that
$\mu(F)\leq\mu\big(B(x,2C\,|X|\cdot|F|)\big)\ll|F|^{\delta},$
where the implied constant depends only on $X$, $C$, and $\mu$.
Next, the second inequality in (2.1) implies that $f$ is Lipschitz with Lipschitz constant at most $C|F|^{-1}$. Hence,
$1=\mu(X)\asymp\mathcal{H}^{\delta}(X)=\mathcal{H}^{\delta}(f(F))\leq(C|F|^{-1})^{\delta}\mathcal{H}^{\delta}(F),$
where the last inequality follows from [17, Proposition 3.1]. Since the $\delta$-Ahlfors regular measure $\mu$ is comparable to $\mathcal{H}^{s}|_{X}$, we have
$\mu(F)\asymp\mathcal{H}^{\delta}|_{X}(F)=\mathcal{H}^{\delta}(F)\geq C^{-\delta}|F|^{\delta},$
as desired. ∎
Proposition 2.6 together with Lemma 2.5 enables us to obtain the following result.
Lemma 2.7.
Let $\mathcal{F}=\{F_{n}\}$ be a sequence of closed sets in $X$ such that the following properties hold:
  1. (a)
    there exists a constant $C\geq 1$ such that for every $F_{n}\in\mathcal{F}$, there exists a bijection $f_{n}:F_{n}\to X$ satisfying
    $C^{-1}\dfrac{|x-y|}{|F_{n}|}\leq|f_{n}(x)-f_{n}(y)|\leq C\dfrac{|x-y|}{|F_{n}|}\text{\qquad for all }x,y\in F_{n};$
  2. (b)
    $\mu\Big(\limsup\limits_{n\to\infty}F_{n}\Big)=1$ and $\lim\limits_{n\to\infty}|F_{n}|=0.$
@@
    If $F_{n}$ and $F_{m}$ are distinct, then $\mathrm{dist}(F_{n},F_{m})\geq\max\{|F_{n}|,|F_{m}|\}$, where the distance between sets is defined by
@@
    $\mu\big(\bigcup_{F_{n}\in\mathcal{F}_{B}}F_{n}\big)\gg\mu(B),$ where the implied constant does not depend on $B$.
@@
If follows from Proposition 2.6 that item (a) of the lemma implies that
@@
where $x_{n}\in F_{n}$ and $c\geq 1$ is a constant depend only on $X$, $C$ and $\mu$. By the full measure assumption in item (b), it is clear that
$\mu\Big(\limsup_{n\to\infty}B(x_{n},c|F_{n}|)\Big)=1.$
This together with another assmuption $\lim_{n\to\infty}|F_{n}|=0$ in item (b), we have that Lemma 2.5 is applicable to the sequence of balls $\{B(x_{n},c|F_{n}|)\}$. By that lemma, we obtain a finite collection $K_{B}$ satisfying items (1)–(3) in Lemma 2.5. Let
@@
By the first inclusion of (2.2), item (1) in the lemma follows immediately from Lemma 2.5 (1) since
$F_{n}\subset B(x_{n},c|F_{n}|)\subset B.$
For item (2), suppose that $F_{n},F_{m}\in\mathcal{F}_{B}$ are distinct. Then, by definition, the same is true for $B(x_{n},c|F_{n}|),B(x_{m},c|F_{m}|)\in K_{B}$. It follows from Lemma 2.5 (2) that
@@
$\mathrm{dist}(F_{n},F_{m})\geq\mathrm{dist}\big(B(x_{n},c|F_{n}|),B(x_{m},c|F_{m}|)\big)\geq\max\{|F_{n}|,|F_{m}|\}.$
For item (3), since $F_{n}$ are pairwise disjoint, we have
[…]
@@
Let $\{E_{n}\}$ be as given in Theorem 1.4. In this section, our goal is to establish that $\limsup E_{n}$ has the large intersection property. Specifically, we will show that for any $0<t<s$,
(3.1) $\limsup_{\ell\to\infty}\mathcal{H}_{\infty}^{t}\bigg(\bigcup_{k=\ell}^{\infty}E_{k}\cap B\bigg)\gg\mu(B)\quad\text{holds for any ball $B\subset X$},$
@@
Remark 7.
Although our proof relies on the original ‘balls to open sets’ mass transference principle (that is, Theorem 2.3 or Theorem 1.4 in the case $\mu=\nu$), it does not immediately imply Theorem 1.4 itself, as can be seen from our proof of that theorem. The reason we employ the original ‘balls to open sets’ mass transference principle is that, as observed in our previous paper [22], this principle has the potential to serve as a general framework for determining the Hausdorff measure/dimension of $\limsup$ sets in a relatively simple manner. Quite surprisingly, this principle continues to be applicable in the present new setting.
@@ 3.1. Construction of a subset of $\bigcup_{k=\ell}^{\infty}E_{k}\cap B$
For the given ball $B$, let $\mathcal{F}_{B}\subset\mathcal{F}$ be the finite subcollection of closed sets obtained from Lemma 2.7. Recall from Definition 1.2 that for any $F_{i}\subset\mathcal{F}$,
@@
where $f_{i}:F_{i}\to X$ is a bijection. Let $F_{i}\in\mathcal{F}_{B}$. Since $f_{i}$ is a bijection, we can estimate the measure of the intersection $B(x,\rho)\cap F_{i}$ for $x\in F_{i}$ and $0<\rho<|F_{i}|$ as follows:
@@
[…]
@@
Then, for any $0<\rho<|F_{i}|/(CN)$ (or equivalently $0<C\rho/|F_{i}|<1/N$),
[…]
@@
Since $\nu(\limsup B(x_{n},r_{n}))=1$ (by Theorem 1.4 (2)), we have $f_{i}^{-1}(E_{\nu}^{N,\varepsilon})=E_{\nu,F_{i}}^{N,\varepsilon}$ except for a set of zero $\nu$-measure. Therefore,
[…]
@@
The above inclusion $B(x_{n_{z}},r_{n_{z}})\subset B$ is possible since $E_{\nu,F_{i}}^{N,\varepsilon}\subset F_{i}\subset B$ and $B$ is open. Set $L_{n_{z}}:=B(z,5r_{n_{z}})$. Recall from Theorem 1.4 that $E_{n}\subset B_{n}$. Then, we have
@@
By item (2) above, there exists some $1\leq k_{i}\leq Q_{d}$ such that the corresponding collection $\mathcal{B}_{k_{i}}(F_{i})$ satisfies
[…]
Since the balls in $\mathcal{B}_{k_{i}}(F_{i})$ are pairwise disjoint (see item (1) above), we may further extract a finite subcollection $\mathcal{B}(F_{i})\subset\mathcal{B}_{k_{i}}(F_{i})$ such that
@@
(1) It follows from immediately from Lemma 2.7.
@@
Remark 8.
Let us explain why it is necessary to go to such lengths to construct $A$ defined in (3.11). First, suppose, for the moment, that $\nu=\mu$ is the $\delta$-Ahlfors regular measure. Then, any ball $B(x,r)\subset F_{i}$ ‘cannot contain too much’ sets $L_{n_{z}}=B(z,5r_{n_{z}})$, as the $\nu$-measure of $B(x,r)$ and $L_{n_{z}}$ is proportional to the $\delta$-power of their respective radii. This essentially means that the sets $L_{n_{z}}$, and hence $E_{n_{z}}$, are well-separated. Therefore, one could expect that the Hausdorff content of $A$, and hence the Hausdorff dimension of $\limsup E_{n}$, is as large as anticipated.
However, the situation becomes subtle when $\nu$ is singular with respect to $\mu$, and we no longer have an effective and uniform estimate for $\nu(B(x,r))$. In such cases, it could happen that $\nu(B(x,r))=r^{\alpha}$, with $\alpha$ much smaller than $\operatorname{\overline{\dim}_{H}}\nu$, and consequently, $B(x,r)$ would ‘contain too many’ sets $L_{n_{z}}$, which would, in turn, lead to $B(x,r)$ containing too many sets $E_{n_{z}}$. In other words, the mass of $A$ may concentrate in a small part of $A$. As a result, the Hausdorff content of $A$ would be smaller than we expected, and so would the Hausdorff dimension of $\limsup E_{n}$.
The notion of quasi-self-conformality of $\nu$ is designed to avoid this situation. As can be seen in the construction of the set $E_{\nu,F_{i}}^{N,\varepsilon}$, the ‘conformal map’ $f_{i}$ preserves some local structure of $\nu$, thus allowing us to focus on the set of points $x\in F_{i}$ that satisfy the desired estimate:
$\frac{\nu(B(x,\rho)\cap F_{i})}{\nu(F_{i})}\ll\bigg(\frac{\rho}{|F_{i}|}\bigg)^{\operatorname{\overline{\dim}_{H}}\nu-\varepsilon}.$